# Some mathematical issues regarding a new approach towards quantum foundation

### v3

@@ 6 Some final remarks
As discussed in Helland (2024a,c), there have been proposed other possible foundations of quantum theory, and my approach should be compared to these. I will claim that the postulates stated here, and also in Helland (2024b,c) are simpler than most proposals in the literature, but detailed arguments behind such a claim are beyond the scope of this article, which has concentrated on the rather simple mathematics behind my approach.
@@
A limitation of my approach is that I do not assume the full validity of the superposition principle. I limit the concept of state vectors to Hilbert space vectors which are eigenvectors of some physically meaningful operators. These can be identified by questions of the form ‘What is $\theta$? /What will $\theta$ be if we measure it?’ for some accessible variable $\theta$, together with sharp answers of the form $\theta=u$. For some such questions, answers of the type ‘We don’t know’ are allowed. Thinking in this way, give simple explanations for paradoxes like Schrödinger’s cat and Wigner’s friend, see Helland (2023c).
The postulates of this article generalize and at the same time simplify the postulates of Helland (2024c), where the symmetry conditions and the question of when a symmetric operator was self-adjoint, were not taken into account. On the other hand, in Helland (2024c) conditions for the validity of the Born formula were discussed. A more thorough discussion of the Born formula is given in Helland (2024b).

### v4

@@ 6 Some final remarks
As discussed in Helland (2024a,c), other possible foundations of quantum theory have been proposed, and my approach should be compared to these. I will claim that the postulates stated here, and also in Helland (2024b,c), are simpler than most proposals in the literature, but detailed arguments behind such a claim are beyond the scope of this article, which has concentrated on the rather simple mathematics behind my approach.
@@
A limitation of my approach is that I do not assume the full validity of the superposition principle. I limit the concept of state vectors to Hilbert space vectors which are eigenvectors of some physically meaningful operators. These can be identified by questions of the form ‘What is $\theta$? /What will $\theta$ be if we measure it?’ for some accessible variable $\theta$, together with sharp answers of the form $\theta=u$. For some such questions, answers of the type ‘We don’t know’ are allowed. Thinking in this way, gives simple explanations for paradoxes like Schrödinger’s cat and Wigner’s friend; see Helland (2023c).
The postulates of this article generalise and at the same time simplify the postulates of Helland (2024c), where the symmetry conditions and the question of when a symmetric operator was self-adjoint were not taken into account. On the other hand, in Helland (2024c), conditions for the validity of the Born formula were discussed. A more thorough discussion of the Born formula is given in Helland (2024b).